OpenStudy (anonymous):

Solve. (x-15)-(2x-11)=3

OpenStudy (anonymous):

x-15-2x+11=3 (distribute the negative) -x-4=3 (add like terms) -x=7 (add 4 to both sides) x=-7 (multiply by negative 1)

OpenStudy (anonymous):

Thank you!